OpenStudy (anonymous):

Find the derivative from the left at x = 2. f(x)=sqrt(4-x^2)

OpenStudy (anonymous):

have you got to the chain rule yet?

OpenStudy (anonymous):

Not yet

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

can't solve without chain rule?

OpenStudy (anonymous):

so you have to compute \[\lim_{x\to 2^-}\frac{\sqrt{4-x^2}}{x-2}\] i guess

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt{4-x^2}\) to get \[\frac{4-x^2}{(x-2)\sqrt{4-x^2}}\] then factor and cancel

OpenStudy (anonymous):

the denominator is still equal to 0, as I cancel (x-2)

OpenStudy (anonymous):

What rules do you get to use?

OpenStudy (anonymous):

Just the definition of a limit?

OpenStudy (anonymous):

\[ \Large \lim_{h\to 0^-}\frac{\sqrt{4-(x+h)^2}-\sqrt{4-x^2}}{h} \]